a pine tree casts a shadow 9 feet long. At the same time, a rod measuring 4 feet casts a shadow 1.5 feet long. how tall is the pine tree ?
Answer by rfer(16322)

x/9=4/1.5
1.5x=36
x=36/1.5
x=24 ft